# Quantifying Memorization and Retriever Performance in Retrieval-Augmented Vision-Language Models

### v1

@@ 8 Conclusion
We demonstrate that retrieval-augmented VLMs have improved performance over general-purpose VLMs, with comparable memorization rates. However, there is still a substantial performance gap between finetuned and baseline QA models. By introducing UCR and PPR, we provide concrete measures of how retrieval mitigates memorization. This analysis outlines the interplay between parametric knowledge and external retrieval. Our measures are validated by the fact that they reveal this well-known tradeoff between memorization and generalization. Our work provides a foundation for future research aimed at refining retrieval mechanisms and ensuring that external sources effectively complement the parametric knowledge of VLMs.
9 Limitations
Our measures are based upon the assumption that incorrectly retrieved sources should result in incorrect answers from the VQA model. They are proxy measures premised upon the assumption that incorrectly retrieved sources do not entail the correct answer. Future work to build memorization metrics not subject to this assumption are warranted.
Given that there is a wealth of research on quantifying the parametric effect using unimodal QA benchmarks, we choose to look only at VQA. That being said, the evaluation metrics we present are general and are also relevant to the measurement of the parametric effect in unimodal settings. We leave this application to future work.
As our memorization metrics require that the task be designed in a two-part retrieval and VQA process, this leaves WebQA as the only valid VQA task to evaluate performance on. Note that we do validate MH-VoLTA performance on VQAv2 and NLVR2 (see Table 3 in the appendix), as this novel QA model should validated independently of external retrieval systems. As the other QA models have been validated in prior works, we do not measure their performance on VQAv2 or NLVR2.
@@ A.1 Model Selection Results
In addition to WebQA, we evaluate models on two fixed-input VQA datasets—VQAv2 (Goyal et al., 2017), a multi-class, single-image VQA dataset, and (Suhr et al., 2018), a binary classification, two-image VQA dataset. These datasets are well-suited to VoLTA classifier architecture. In particular, question categories in VQAv2, along with the associated answer-domains, match well with WebQA, with a substantial portion of both datasets focusing on color, shape, number, and yes/no questions.
@@ A.3 Performance by Question Category
We report the mean accuracy per question category for Multihop-VoLTA in 5(a) using source retrieval oracles. We find that performance is dependent upon the level of training data available, with the shape category having the least number of samples in the dataset. Question counts per category are as follows; Yes/No (n = 7,320), color (n = 1,830), number (n = 2,118), shape (n = 565). The similarity in results across different question categories reinforces the reliability and stability of our model’s performance. For a breakdown of labels per question category, see subsection A.6.
@@ A.4 GPT-4o Retrieval Prompt
A.5 Question Complexity Analysis Metrics
@@
A.6 Question Category Domain Lists
@@
A.7 Baseline Models
VLP
$A^{l}=softmax(\frac{Q^{T}K}{\sqrt{d}}+M)V^{T}$ (13)
GIT
BLIP-2
GPT-3.5 Turbo

### v2

@@ 8 Conclusion
We demonstrate that retrieval-augmented VLMs have improved performance over general-purpose VLMs, with comparable parametric response rates. However, there is still a substantial performance gap between finetuned and baseline QA models. By introducing UCR and PPR, we provide concrete measures of how incrementally improving retrieval systems mitigates parametric responses. This analysis outlines the interplay between parametric knowledge and external retrieval, highlighting well-known tradeoff between memorization and generalization in the multimodal setting. Finally, we demonstrate that current VLMs have higher parameteric response rates when reasoning over image sources rather than text sources. Our work provides a foundation for future research aimed at refining retrieval mechanisms and ensuring that external sources effectively complement the parametric knowledge of VLMs.
@@ A.1 Model Selection Results
As our memorization metrics require that the task be designed in a two-part retrieval and VQA process, this leaves WebQA as the only valid VQA task to evaluate performance on. Here, independently of any external retrieval system, we validate MH-VoLTA performance on two fixed-input VQA datasets (see Table 3). As the other baseline models have been validated in prior works, we do not measure their performance on VQAv2 or NLVR2 again.
We evaluate models VQAv2 (Goyal et al., 2017), a multi-class, single-image VQA dataset, and (Suhr et al., 2018), a binary classification, two-image VQA dataset. These datasets are well-suited to VoLTA classifier architecture. In particular, question categories in VQAv2, along with the associated answer-domains, match well with WebQA, with a substantial portion of both datasets focusing on color, shape, number, and yes/no questions.
@@ A.3 Performance by Question Category
We report the mean accuracy per question category for Multihop-VoLTA in 5(a) using source retrieval oracles. We find that performance is dependent upon the level of training data available, with the shape category having the least number of samples in the dataset. Question counts per category are as follows; Yes/No (n = 7,320), color (n = 1,830), number (n = 2,118), shape (n = 565). The similarity in results across different question categories reinforces the reliability and stability of our model’s performance. For a breakdown of labels per question category, see Figure 8.
@@ A.4 GPT-4o Retrieval Prompt
A.5 Robustness Checks: Question Complexity
Complexity Metrics
To identify correlations between the complexity of the question, retrieval recall, and QA accuracy, we apply three separate measures to the input questions; Word Count, Flesch-Kincaid Grade Level (Flesch, 2007), and Gunning-Fog Index (Gunning, 1952).
@@ A.5 Robustness Checks: Question Complexity
Complexity Analysis
We observe and report interesting relationships between query complexity and retrieval and QA performance. We find that the accuracy of the in-context GPT-4o retriever is related to question complexity (Figure 6). The more complex the question in terms of word count, Flesch-Kincaid Grade, or Gunning Fog Index, the lower the QA performance (see Figure 6). In contrast, increasing query complexity improves GPT-4o’s retrieval ability, where the additional complexity provides information on source relevancy. However, this relationship does not hold for the finetuned UniVL-DR retriever, where question complexity has little effect on retrieval recall or QA accuracy (Figure 7). As such, systems that rely on "in-context" retrieval using GPT-4o are limited by query complexity, but approaches that utilize finetuned retrievers are not.
We note that the opposing relationship between retrieval and QA performance is contrary to the finding that the QA task is heavily dependent on retrieval performance (2(b)). The impact of query complexity on task performance is strong enough to overcome this general principle.
A.6 Baseline Models
VLP
$A^{l}=softmax(\frac{Q^{T}K}{\sqrt{d}}+M)V^{T}$ (13)
GIT
BLIP-2
GPT-3.5 Turbo
@@ A.6 Baseline Models
Figure 8: WebQA keyword lists per question category.